a)Mary invested $600 at the Royal Bank at an interest rate of 3.6% per year.How much money will she have after 12 years?
b)The half-life of radium is 1400 years. If there was 800 grams of radium, how much will remain after 5000 years?

Answer by stanbon(75887)

a)Mary invested $600 at the Royal Bank at an interest rate of 3.6% per year.How much money will she have after 12 years?
Interest = (pricipal)*(rate)*(time)
I = 600*0.036*12 = $259.20
---------------------
b)The half-life of radium is 1400 years. If there was 800 grams of radium, how much will remain after 5000 years?
Formula: A(t)= A(0)(1/2)^t/h, where h is the half-life
Your Problem:
A(5000) = 800(1/2)^(5000/1400)
A(5000) = 800*0.0841
A(5000) = 67.30 grams
===============
Cheers,
Stan H.